# Stability of discrete-symmetry flocks: sandwich state, traveling domains and motility-induced pinning

### v1

@@ III.2 Numerical simulations with droplet excitation
We now undertake a detailed analysis of the 4-state APM using Monte Carlo simulations, which constitute the core of our study and enable a thorough examination of the stability of the polar-ordered liquid phase. The stability of this phase is assessed on two-dimensional square and rectangular lattices by introducing either a counter-propagating or a transversely propagating liquid droplet into the high-density ordered liquid, in line with the coarse-grained hydrodynamic framework. A schematic diagram of this arrangement is shown in Fig. 2(i–ii). The simulation protocol follows Ref. Benvegnen et al. (2023), where the initial liquid configuration is constructed with particles of $\sigma=3$ (blue) state and thermalized up to a time $t_{\rm th}=150$ to allow the system to adapt to the changed environment. At $t=0$, a circular region of radius $r_{d}$, centered at $(L_{x}/2,L_{y}/2)$, is chosen, and an additional $\Delta N=(\rho_{0}^{d}-\rho_{0})\pi r_{d}^{2}$ particles are inserted to form a high-density droplet ($\rho_{0}^{d}>\rho_{0}$). Here, $\rho_{0}^{d}$ and $r_{d}$ denote the droplet’s density and radius, respectively. The internal state of all particles in the droplet is then set to $\sigma=1$ (red) for a counter-propagating droplet, or to $\sigma=2$ (green) or $\sigma=4$ (black) for a transversely propagating droplet. Simulations are carried out across a range of control parameters: the diffusion constant $D$, the inverse temperature $\beta=1/T$ controlling noise, and the self-propulsion strength $\epsilon$, which sets the effective particle speed. The system is evolved using a random-sequential-update Monte Carlo algorithm until a stationary state is reached at $t_{\rm eq}\sim 10^{3}$$10^{5}$, after which ensemble-averaged measurements are recorded.
@@
Fig. 2 illustrates the time evolution and resulting steady states following artificial droplet excitations at different self-propulsion velocities, deep inside the liquid phase ($\beta=1$, $\rho_{0}=10$). As shown in Fig. 2(a–f), a small counter-propagating droplet does not destabilize the polar ordered phase at a low bias of $\epsilon=0.3$ [Fig. 2(a–b)], but it disrupts the liquid phase at intermediate and high biases, $\epsilon=0.9$ and $\epsilon=2.7$ [Fig. 2(c–f)], thereby confirming the metastable character of the polar liquid phase in the 4-state APM. For $\epsilon=0.3$ [Fig. 2(a–b)], the directional hopping rate is low, while the non-directional hopping rates, although smaller, remain appreciable [see Eq. (4)]. As a consequence, the droplet dissolves gradually without perturbing the original ordered phase composed of $\sigma=3$ particles. At $\epsilon=0.9$ [Fig. 2(c–d)], we observe that the droplet affects the ordered background similarly to the behavior seen in the AIM Benvegnen et al. (2023), expanding at the cost of the ordered region and leaving behind a comet-shaped trail as it advances along its preferred direction (see Appendix A for the time evolution of such a counter-propagating droplet at low and high self-propulsion velocities). This growth results from fluctuations in orientation at the droplet boundary, where blue $\sigma=3$ particles convert into red $\sigma=1$ particles, fueling further expansion. As the transformation proceeds, the original polar phase (blue, $\sigma=3$) is entirely replaced by a new ordered phase associated with the droplet (red, $\sigma=1$), as seen in Fig. 2(d). In contrast, the limited transverse hopping along the ±$\pm y$-direction at higher propulsion $\epsilon=2.7$ prevents significant transverse expansion of the droplet [Fig. 2(e–f)]. As discussed in the context of Fig. 1, this constraint leads to the development of a partially reversed mixed sandwich configuration [Fig. 2(f)], in which regions of the original ordered phase (blue) coexist with a red lane formed by the longitudinal motion of the droplet. This sandwich structure constitutes a stable steady-state that was not reported in previous studies involving droplet excitations Codina et al. (2022); Benvegnen et al. (2023), although such a configuration can emerge for reduced transverse diffusion (see Appendix B for more detail). Importantly, the lane’s final position in Fig. 2(f) depends on the initial placement of the droplet, indicating that multiple stable sandwich steady-states may arise. This structure can also emerge spontaneously in the APM for large bias $\epsilon$, even in the absence of any initial droplet perturbation, when starting from random configurations.
@@
Although Fig. 2(d) illustrates a full reversal of the initial liquid phase triggered by a counter-propagating droplet, this apparent reversal is in fact a finite-size effect. Let $w_{\rm sw}$ be the sandwich width for the infinite system, then one observes complete reversal in a finite system when $L_{y}<w_{\rm sw}$ (keeping $L_{x}$ fixed). We, therefore, conclude that deep inside the liquid phase (low noise) and in the thermodynamic limit, the sandwich state becomes the unique steady-state outcome following a counter-propagating droplet perturbation. See Appendix C for a detailed discussion of how spatial anisotropy governs the stability of the APM sandwich state. In what follows, we provide a quantitative analysis of how $L_{x}$, $L_{y}$, and $\epsilon$ influence the emergence of this configuration.
Figure 3: (color online) Phase reversal due to counter-propagating droplet. (a) Average relative width ($w_{sw}/L_{y}$) of the sandwich state versus aspect ratio of the simulation box $(L_{x}/L_{y})$ for fixed $r_{d}=10$ and $\epsilon=0.9$. $w_{sw}/L_{y}=1$ corresponds to a full reversal of the initial ordered state. (b) $w_{sw}/L_{y}$ as a function of relative droplet size $(2r_{d}/L_{y})$ obtained for fixed $r_{d}=6$ and $\epsilon=0.9$. (c) $w_{sw}/L_{y}$ as a function of $\epsilon$ for counter-propagating droplet with fixed $L_{y}=50$ and $r_{d}=10$. A movie (movie03) showing the sandwich formation for various $\epsilon$ can be found at Ref. Chatterjee et al. (2025). Parameters: $D=1$, $\beta=1$, $\rho_{0}=10$, $\rho_{0}^{d}=5\rho_{0}$ and (a) $L_{y}=50$, (b–c) $L_{x}=200$.
@@
Fig. 3 illustrates how spatial anisotropy and self-propulsion strength influence the reversal dynamics of the flocking phase under the action of counter-propagating droplets. In Fig. 3(a), the relative width of the sandwich region, $w_{\rm sw}/L_{y}$, is plotted against the aspect ratio of the simulation domain at a fixed transverse size $L_{y}=50$. As $L_{x}$ increases, the system undergoes a transition from a sandwich configuration $(w_{\rm sw}/L_{y}<1)$ to a full reversal $(w_{\rm sw}/L_{y}=1)$ of the original ordered phase. This behavior arises because a larger longitudinal size allows the droplet more time to grow and move before its front encounters its trailing end, enabling it to span the entire system vertically. In contrast, for smaller $L_{x}$, the growth of the droplet front is interrupted by its own trajectory, leading to a persistent sandwich state. Fig. 3(b) shows that, for fixed droplet radius $r_{d}=6$, the sandwich width $w_{\rm sw}/L_{y}$ exhibits a discontinuous change as a function of $L_{y}$. Complete reversal is favored when $L_{y}$ is close to the droplet diameter, but becomes impossible as $L_{y}$ grows significantly beyond it. In this regime, the width of the sandwich region $w_{\rm sw}$ becomes independent of $L_{y}$, leading to a linear increase of $w_{\rm sw}/L_{y}$ with $1/L_{y}$. Thus, full reversal is favored only at small transverse sizes. For large self-propulsion $\epsilon$, this linear trend persists even for smaller $L_{y}$, and full reversal occurs only when $L_{y}\approx 2r_{d}$. In Fig. 3(c), $w_{\rm sw}/L_{y}$ is plotted against $\epsilon$ for $L_{y}=50$. At low $\epsilon$, transverse diffusion is significant, resulting in $w_{\rm sw}>2r_{d}$. As $\epsilon$ increases, transverse hopping is suppressed, and $w_{\rm sw}/L_{y}$ decreases gradually until around $\epsilon\sim 1.4$, the point at which the system transitions from transverse to longitudinal motion Chatterjee et al. (2020); Mangeat et al. (2020). Beyond $\epsilon\gtrsim 2$, transverse motion becomes negligible, and the sandwich width saturates at approximately $w_{\rm sw}\simeq 2r_{d}-1$. Across this range, the sandwich state emerges as the generic steady state for counter-propagating droplet excitations. This trend is similar to the one obtained from the hydrodynamic theory shown in Fig. 1(f).
@@
Results from the APM hydrodynamic theory [Sec. III.1] indicate that the system exhibits qualitatively different responses to counter-propagating droplet excitations depending on the noise strength [Fig. 1(a–b)]. At low noise, the droplet stabilizes into a sandwich configuration, whereas at high noise, it can fully destabilize the background liquid phase. We have already discussed in length the low-noise regime deep inside the liquid phase, where the sandwich state is the dominant and stable outcome. Now, we return to the high-noise regime to demonstrate that for stronger noise strength $(\beta\leq 0.82)$, the stability of the sandwich state weakens, and a counter-propagating droplet can completely reverse the initial liquid phase, in line with the hydrodynamic prediction (see Fig. 1(b), Appendix D and movie04 in Ref. Chatterjee et al. (2025) for further details). In Fig. 4, we systematically explore the response of the system with counter-propagating droplet excitation by constructing comprehensive phase diagrams across the full range of control parameters ($\beta$, $\epsilon$, and $\rho_{0}$). The $\beta-\epsilon$ diagram [Fig. 4(a)] is shown for a very high density $(\rho_{0}=10)$ and dominated by the sandwich configuration (I) when noise is small and the system is deep inside the liquid phase. However, as noise increases, the system undergoes complete reversal (II) for intermediate values of self-propulsion, where the counter-propagating droplet progressively displaces and replaces the initial liquid phase. Both the $\beta-\rho_{0}$ and $\epsilon-\rho_{0}$ diagrams [Figs. 4(b–c)] reveal a rich variety of steady states as the system moves away from the liquid phase, either by increasing noise or decreasing density. These include a high-density band of the droplet state forming a liquid-gas coexistence regime after destroying the initial liquid (III), a sandwich configuration in the coexistence regime (IV), a large liquid blob of the droplet state (V), pure gas phase (VI), and longitudinal lane formation at higher $\epsilon$ (VII). While the phase boundaries in the diagrams are obtained for a finite system size ($L=200$) and may shift slightly with increasing system size, the overall structure and classification of steady states remain robust. To demonstrate that these phases are not finite-size artifacts, we provide representative snapshots of the steady-states for a larger system size ($L=800$), confirming the persistence and stability of the identified states in the thermodynamic limit.
@@
In the case of a droplet excitation moving transversely, the resulting steady-state configuration also depends sensitively on the particle self-propulsion velocity. At intermediate values of $\epsilon$, the system predominantly exhibits a morphology where two distinct high-density clusters coexist, each maintaining persistent motion in perpendicular directions—one representing the inserted droplet and the other corresponding to the preexisting liquid phase (see Fig. 1(d), Appendix D and movie05 at Ref. Chatterjee et al. (2025) for $0.9\leq\epsilon\leq 1.5$). These clusters, though they frequently collide, do not merge into a single homogeneous phase, highlighting a dynamical mismatch in their collective movement. At higher self-propulsion strengths, however, the transverse droplet fully destabilizes the initial liquid, and the system transitions into a final state comprised entirely of particles in the droplet’s spin state, or into a mixed phase containing both spin types that support transverse motion with respect to the direction of the original liquid flow (see movie05 at Ref.Chatterjee et al. (2025) for $\epsilon\geq 2.4$).
@@
Thus far, we have shown that droplet insertion into the liquid phase of the APM can destabilize it, resulting either in a sandwich state or a complete reversal of the ordered phase. We now provide a quantitative analysis of how the size and density of the droplet influence this reversal. Figs. 5(a,c) display the probability of sandwich formation, $P_{r}$, for counter-propagating droplets as functions of droplet radius $(r_{d})$ and droplet density $(\rho_{0}^{d})$, respectively. In Fig. 5(a), the dashed line shows a fit using a hyperbolic tangent function Benvegnen et al. (2023), which is used to identify the critical radius, $r_{c}$, marking the transition from the initial ordered phase to the sandwich state triggered by a counter-propagating droplet. The self-propulsion velocity, $\epsilon$, also significantly affects the phase stability. Fig. 5(b) shows how the critical radius $r_{c}$ varies with $\epsilon$. At higher bias, smaller droplets are sufficient to generate a sandwich state, due to enhanced hopping in the preferred direction and suppressed transverse movement. As $\epsilon$ decreases and transverse diffusion becomes more prominent, a larger $r_{c}$ is needed to achieve the transition. The phase diagram in Fig. 5(d), plotted in the $(r_{d},\rho_{0}^{d})$ plane, indicates that the sandwich state remains the dominant steady-state configuration with counter-propagating droplet excitation. This phase diagram resembles the one obtained from the hydrodynamic theory shown in Fig. 1(g).
@@
Fig. 6, which corresponds to the case of transverse droplet propagation, is analogous to Fig. 5 for counter-propagating droplets. It presents how transverse motion relative to the homogeneous liquid phase affects phase reversal. An increase in the initial droplet density $\rho_{0}^{d}$ tends to stabilize the droplet, increasing the likelihood of a complete reversal of the ordered phase. Similarly, a larger droplet radius $r_{d}$ enhances the probability of reversal, indicating that both parameters play key roles in determining the stability of the phase. Nonetheless, compared to counter-propagating droplets, a stronger perturbation is needed to achieve complete reversal in the transverse case [Figs. 6(a,c)]. A comparison between Figs. 6(a,c) and Figs. 5(a,c) shows that counter-propagating droplets can destabilize the ordered phase more efficiently, requiring smaller $r_{d}$ and lower $\rho_{0}^{d}$ than their transversely propagating counterparts. In the case of transverse propagation, the critical radius $r_{c}$ necessary for full reversal increases sharply as $\epsilon$ decreases, unlike in the counter-propagating scenario [Fig. 6(b)]. At small $\epsilon$, enhanced non-biased hopping leads to greater diffusion at the droplet boundary, causing the transversely moving droplet to dissipate. As a result, a significantly larger droplet is needed to induce full reversal when $\epsilon$ is low. The stability diagram in Fig. 6(d), plotted in the $(\rho_{0}^{d},r_{d})$ plane, further characterizes the behavior for transverse droplet propagation. The fully reversed state, when it occurs, features two high-density clusters moving in perpendicular directions, corresponding to the droplet and the initial liquid, as described in Appendix D. This phase diagram is similar to the one obtained from the hydrodynamic theory shown in Fig. 1(h). All these findings with a droplet excitation are then consistent with the results obtained from the hydrodynamic analysis presented in Sec. III.1.
@@
So far, we have reported results for artificial droplet excitation in the APM at a diffusion constant $D=1$, consistent with earlier studies of the model Chatterjee et al. (2020); Mangeat et al. (2020), which primarily focused on steady states at this representative value. However, recent investigations Benvegnen et al. (2023); Woo and Noh (2024); Mangeat et al. (2025) have demonstrated that lower diffusion can substantially influence the emergent collective dynamics in discrete flocking systems. This motivates a more systematic analysis of diffusion’s role in our model, as presented in Fig. 7, which shows four stability diagrams characterizing the droplet-induced behavior of the APM for reduced diffusion. For $\epsilon=1.2$ [Fig. 7(a–b)], the diagrams indicate that achieving either a sandwich state [Fig. 7(a)] or a full reversal [Fig.7(b)] of the initial liquid phase requires higher diffusion as the droplet radius decreases. At this value of $\epsilon$, hopping in non-preferred directions remains appreciable, meaning that only larger droplets can persist when $D$ is small; smaller droplets dissipate over time. As a result, for $D\leq 0.5$, no droplets survive, regardless of their size (up to the maximum tested $r_{d}=10$). Notably, in Fig.7(b), the full reversal state observed at $D\geq 0.7$ and larger $r_{d}$ $(r_{d}\gtrsim 7)$ is marked by orthogonal motion of clusters associated with the droplet and the initial liquid, as detailed in AppendixD.
[…]

### v2

@@ III.2 Numerical simulations with droplet excitation
We now undertake a detailed analysis of the 4-state APM using Monte Carlo simulations, which constitute the core of our study and enable a thorough examination of the stability of the polar-ordered liquid phase. The stability of this phase is assessed on two-dimensional square and rectangular lattices by introducing either a counter-propagating or a transversely propagating liquid droplet into the high-density ordered liquid, in line with the coarse-grained hydrodynamic framework. A schematic diagram of this arrangement is shown in Fig. 2(i–ii). The simulation protocol follows Ref. benvegnen2023meta, where the initial liquid configuration is constructed with particles of $\sigma=3$ (blue) state and thermalized up to a time $t_{\rm th}=150$ to allow the system to adapt to the changed environment. At $t=0$, a circular region of radius $r_{d}$, centered at $(L_{x}/2,L_{y}/2)$, is chosen, and an additional $\Delta N=(\rho_{0}^{d}-\rho_{0})\pi r_{d}^{2}$ particles are inserted to form a high-density droplet ($\rho_{0}^{d}>\rho_{0}$). Here, $\rho_{0}^{d}$ and $r_{d}$ denote the droplet’s density and radius, respectively. The internal state of all particles in the droplet is then set to $\sigma=1$ (red) for a counter-propagating droplet, or to $\sigma=2$ (green) or $\sigma=4$ (black) for a transversely propagating droplet. Simulations are carried out across a range of control parameters: the diffusion constant $D$, the inverse temperature $\beta=1/T$ controlling noise, and the self-propulsion strength $\epsilon$, which sets the effective particle speed. The system is evolved using a random-sequential-update Monte Carlo algorithm until a stationary state is reached at $t_{\rm eq}\sim 10^{3}$$10^{5}$, after which ensemble-averaged measurements are recorded.
@@
Fig. 2 illustrates the time evolution and resulting steady states following artificial droplet excitations at different self-propulsion velocities, deep inside the liquid phase ($\beta=1$, $\rho_{0}=10$). As shown in Fig. 2(a–f), a small counter-propagating droplet does not destabilize the polar ordered phase at a low bias of $\epsilon=0.3$ [Fig. 2(a–b)], but it disrupts the liquid phase at intermediate and high biases, $\epsilon=0.9$ and $\epsilon=2.7$ [Fig. 2(c–f)], thereby confirming the metastable character of the polar liquid phase in the 4-state APM. For $\epsilon=0.3$ [Fig. 2(a–b)], the directional hopping rate is low, while the non-directional hopping rates, although smaller, remain appreciable [see Eq. (4)]. As a consequence, the droplet dissolves gradually without perturbing the original ordered phase composed of $\sigma=3$ particles. At $\epsilon=0.9$ [Fig. 2(c–d)], we observe that the droplet affects the ordered background similarly to the behavior seen in the AIM benvegnen2023meta, expanding at the cost of the ordered region and leaving behind a comet-shaped trail as it advances along its preferred direction (see Appendix A for the time evolution of such a counter-propagating droplet at low and high self-propulsion velocities). This growth results from fluctuations in orientation at the droplet boundary, where blue $\sigma=3$ particles convert into red $\sigma=1$ particles, fueling further expansion. As the transformation proceeds, the original polar phase (blue, $\sigma=3$) is entirely replaced by a new ordered phase associated with the droplet (red, $\sigma=1$), as seen in Fig. 2(d). In contrast, the limited transverse hopping along the $\pm y$-direction at higher propulsion $\epsilon=2.7$ prevents significant transverse expansion of the droplet [Fig. 2(e–f)]. As discussed in the context of Fig. 1, this constraint leads to the development of a partially reversed mixed sandwich configuration [Fig. 2(f)], in which regions of the original ordered phase (blue) coexist with a red lane formed by the longitudinal motion of the droplet. This sandwich structure constitutes a stable steady-state that was not reported in previous studies involving droplet excitations codina2022small; benvegnen2023meta, although such a configuration can emerge for reduced transverse diffusion (see Appendix B for more detail). Importantly, the final position of the lane in Fig. 2(f) depends on the initial placement of the droplet, indicating that multiple stable sandwich steady-states may arise. This structure can also emerge spontaneously in the APM for large bias $\epsilon$, even in the absence of any initial droplet perturbation, when starting from random configurations.
@@
Although Fig. 2(d) illustrates a full reversal of the initial liquid phase triggered by a counter-propagating droplet, this apparent reversal is in fact a finite-size effect. Let $w_{\rm sw}$ be the sandwich width for the infinite system, then one observes a complete reversal in a finite system when $L_{y}<w_{\rm sw}$ (keeping $L_{x}$ fixed). We, therefore, conclude that deep inside the liquid phase (low noise) and in the thermodynamic limit, the sandwich state becomes the unique steady-state outcome following a counter-propagating droplet perturbation. See Appendix C for a detailed discussion of how spatial anisotropy governs the stability of the APM sandwich state. In what follows, we provide a quantitative analysis of how $L_{x}$, $L_{y}$, and $\epsilon$ influence the emergence of this configuration.
Figure 3: (color online) Phase reversal due to counter-propagating droplet. (a) Average relative width ($w_{sw}/L_{y}$) of the sandwich state versus aspect ratio of the simulation box $(L_{x}/L_{y})$ for fixed $r_{d}=10$ and $\epsilon=0.9$. $w_{sw}/L_{y}=1$ corresponds to a full reversal of the initial ordered state. (b) $w_{sw}/L_{y}$ as a function of relative droplet size $(2r_{d}/L_{y})$ obtained for fixed $r_{d}=6$ and $\epsilon=0.9$. (c) $w_{sw}/L_{y}$ as a function of $\epsilon$ for counter-propagating droplet with fixed $L_{y}=50$ and $r_{d}=10$. A movie (movie03) showing the sandwich formation for various $\epsilon$ can be found at Ref. zenodo. Parameters: $D=1$, $\beta=1$, $\rho_{0}=10$, $\rho_{0}^{d}=5\rho_{0}$ and (a) $L_{y}=50$, (b–c) $L_{x}=200$.
@@
Fig. 3 illustrates how spatial anisotropy and self-propulsion strength influence the reversal dynamics of the flocking phase under the action of counter-propagating droplets. In Fig. 3(a), the relative width of the sandwich region, $w_{\rm sw}/L_{y}$, is plotted against the aspect ratio of the simulation domain at a fixed transverse size $L_{y}=50$. As $L_{x}$ increases, the system undergoes a transition from a sandwich configuration $(w_{\rm sw}/L_{y}<1)$ to a full reversal $(w_{\rm sw}/L_{y}=1)$ of the original ordered phase. This behavior arises because a larger longitudinal size allows the droplet more time to grow and move before its front encounters its trailing end, enabling it to span the entire system vertically. In contrast, for smaller $L_{x}$, the growth of the droplet front is interrupted by its own trajectory, leading to a persistent sandwich state. Fig. 3(b) shows that, for fixed droplet radius $r_{d}=6$, the sandwich width $w_{\rm sw}/L_{y}$ exhibits a discontinuous change as a function of $L_{y}$. Complete reversal is favored when $L_{y}$ is close to the droplet diameter, but becomes impossible as $L_{y}$ grows significantly beyond it. In this regime, the width of the sandwich region $w_{\rm sw}$ becomes independent of $L_{y}$, leading to a linear increase of $w_{\rm sw}/L_{y}$ with $1/L_{y}$. Thus, full reversal is favored only at small transverse sizes. For large self-propulsion $\epsilon$, this linear trend persists even for smaller $L_{y}$, and full reversal occurs only when $L_{y}\approx 2r_{d}$. In Fig. 3(c), $w_{\rm sw}/L_{y}$ is plotted against $\epsilon$ for $L_{y}=50$. At low $\epsilon$, transverse diffusion is significant, resulting in $w_{\rm sw}>2r_{d}$. As $\epsilon$ increases, transverse hopping is suppressed, and $w_{\rm sw}/L_{y}$ decreases gradually until around $\epsilon\sim 1.4$, the point at which the system transitions from transverse to longitudinal motion chatterjee2020flocking; mangeat2020flocking. Beyond $\epsilon\gtrsim 2$, transverse motion becomes negligible, and the sandwich width saturates at approximately $w_{\rm sw}\simeq 2r_{d}-1$. Across this range, the sandwich state emerges as the generic steady state for counter-propagating droplet excitations. This trend is similar to the one obtained from the hydrodynamic theory shown in Fig. 1(f).
Additionally, we report in Appendix D that the sandwich state is not density-segregated. The density profile, integrated over the x-coordinate, remains uniform; only the internal polarization changes along the transverse direction. Appendix D also presents a stability diagram of the sandwich state, identifying that this state remains stable only if the initial width is between two thresholds: $w_{\min}<w_{\rm sw}<w_{\max}$, otherwise the system relaxes to a liquid phase of a single state. Therefore, for a given set of control parameters $(D,\beta,\rho_{0})$, one can prepare the system either in the uniform ordered flocking state or in a sandwich state with any width in the interval $\in[w_{\min}(\epsilon),w_{\max}(\epsilon)]$, and all such states are stable. In this sense, the model admits a continuous family of stable sandwich states with distinct widths. A localized droplet perturbation applied to the ordered flocking state just drives the system into one member of this family, selecting a final width within the stability window.
@@
Results from the APM hydrodynamic theory [Sec. III.1] indicate that the system exhibits qualitatively different responses to counter-propagating droplet excitations depending on the noise strength [Fig. 1(a–b)]. At low noise, the droplet stabilizes into a sandwich configuration, whereas at high noise, it can fully destabilize the background liquid phase. We have already discussed in length the low-noise regime deep inside the liquid phase, where the sandwich state is the dominant and stable outcome. Now, we return to the high-noise regime to demonstrate that for stronger noise strength $(\beta\leq 0.82)$, the stability of the sandwich state weakens, and a counter-propagating droplet can completely reverse the initial liquid phase, in line with the hydrodynamic prediction (see Fig. 1(b), Appendix E and movie04 in Ref. zenodo for further details). In Fig. 4, we systematically explore the response of the system with counter-propagating droplet excitation by constructing comprehensive phase diagrams across the full range of control parameters ($\beta$, $\epsilon$, and $\rho_{0}$). The $\beta-\epsilon$ diagram [Fig. 4(a)] is shown for a very high density $(\rho_{0}=10)$ and dominated by the sandwich configuration (I) when noise is small and the system is deep inside the liquid phase. However, as noise increases, the system undergoes complete reversal (II) for intermediate values of self-propulsion, where the counter-propagating droplet progressively displaces and replaces the initial liquid phase. Both the $\beta-\rho_{0}$ and $\epsilon-\rho_{0}$ diagrams [Figs. 4(b–c)] reveal a rich variety of steady states as the system moves away from the liquid phase, either by increasing noise or decreasing density. These include a high-density band of the droplet state forming a liquid-gas coexistence regime after destroying the initial liquid (III), a sandwich configuration in the coexistence regime (IV), a large liquid blob of the droplet state (V), pure gas phase (VI), and longitudinal lane formation at higher $\epsilon$ (VII). While the phase boundaries in the diagrams are obtained for a finite system size ($L=200$) and may shift slightly with increasing system size, the overall structure and classification of steady states remain robust. To demonstrate that these phases are not finite-size artifacts, we provide representative snapshots of the steady-states for a larger system size ($L=800$), confirming the persistence and stability of the identified states in the thermodynamic limit.
@@
In the case of a droplet excitation moving transversely, the resulting steady-state configuration also depends sensitively on the particle self-propulsion velocity. At intermediate values of $\epsilon$, the system predominantly exhibits a morphology where two distinct high-density clusters coexist, each maintaining persistent motion in perpendicular directions—one representing the inserted droplet and the other corresponding to the preexisting liquid phase (see Fig. 1(d), Appendix E and movie05 at Ref. zenodo for $0.9\leq\epsilon\leq 1.5$). These clusters, though they frequently collide, do not merge into a single homogeneous phase, highlighting a dynamical mismatch in their collective movement. At higher self-propulsion strengths, however, the transverse droplet fully destabilizes the initial liquid, and the system transitions into a final state comprised entirely of particles in the droplet’s spin state, or into a mixed phase containing both spin types that support transverse motion with respect to the direction of the original liquid flow (see movie05 at Ref.zenodo for $\epsilon\geq 2.4$).
@@
Thus far, we have shown that droplet insertion into the liquid phase of the APM can destabilize it, resulting either in a sandwich state or a complete reversal of the ordered phase. We now provide a quantitative analysis of how the size and density of the droplet influence this reversal. Figs. 5(a,c) display the probability of sandwich formation, $P_{r}$, for counter-propagating droplets as functions of droplet radius $(r_{d})$ and droplet density $(\rho_{0}^{d})$, respectively. In Fig. 5(a), the dashed line shows a fit using a hyperbolic tangent function benvegnen2023meta, which is used to identify the critical radius, $r_{c}$, marking the transition from the initial ordered phase to the sandwich state triggered by a counter-propagating droplet. The self-propulsion velocity, $\epsilon$, also significantly affects the phase stability. Fig. 5(b) shows how the critical radius $r_{c}$ varies with $\epsilon$. At higher bias, smaller droplets are sufficient to generate a sandwich state, due to enhanced hopping in the preferred direction and suppressed transverse movement. As $\epsilon$ decreases and transverse diffusion becomes more prominent, a larger $r_{c}$ is needed to achieve the transition. The phase diagram in Fig. 5(d), plotted in the $(r_{d},\rho_{0}^{d})$ plane, indicates that the sandwich state remains the dominant steady-state configuration with counter-propagating droplet excitation. This phase diagram resembles the one obtained from the hydrodynamic theory shown in Fig. 1(g).
@@
Fig. 6, which corresponds to the case of transverse droplet propagation, is analogous to Fig. 5 for counter-propagating droplets. It presents how transverse motion relative to the homogeneous liquid phase affects phase reversal. An increase in the initial droplet density $\rho_{0}^{d}$ tends to stabilize the droplet, increasing the likelihood of a complete reversal of the ordered phase. Similarly, a larger droplet radius $r_{d}$ enhances the probability of reversal, indicating that both parameters play key roles in determining the stability of the phase. Nonetheless, compared to counter-propagating droplets, a stronger perturbation is needed to achieve complete reversal in the transverse case [Figs. 6(a,c)]. A comparison between Figs. 6(a,c) and Figs. 5(a,c) shows that counter-propagating droplets can destabilize the ordered phase more efficiently, requiring smaller $r_{d}$ and lower $\rho_{0}^{d}$ than their transversely propagating counterparts. In the case of transverse propagation, the critical radius $r_{c}$ necessary for full reversal increases sharply as $\epsilon$ decreases, unlike in the counter-propagating scenario [Fig. 6(b)]. At small $\epsilon$, enhanced non-biased hopping leads to greater diffusion at the droplet boundary, causing the transversely moving droplet to dissipate. As a result, a significantly larger droplet is needed to induce full reversal when $\epsilon$ is low. The stability diagram in Fig. 6(d), plotted in the $(\rho_{0}^{d},r_{d})$ plane, further characterizes the behavior for transverse droplet propagation. The fully reversed state, when it occurs, features two high-density clusters moving in perpendicular directions, corresponding to the droplet and the initial liquid, as described in Appendix E. This phase diagram is similar to the one obtained from the hydrodynamic theory shown in Fig. 1(h). All these findings with a droplet excitation are then consistent with the results obtained from the hydrodynamic analysis presented in Sec. III.1.
@@
So far, we have reported results for artificial droplet excitation in the APM at a diffusion constant $D=1$, consistent with earlier studies of the model chatterjee2020flocking; mangeat2020flocking, which primarily focused on steady states at this representative value. However, recent investigations benvegnen2023meta; jd2024MIP; TSAIM have demonstrated that lower diffusion can substantially influence the emergent collective dynamics in discrete flocking systems. This motivates a more systematic analysis of diffusion’s role in our model, as presented in Fig. 7, which shows four stability diagrams characterizing the droplet-induced behavior of the APM for reduced diffusion. For $\epsilon=1.2$ [Fig. 7(a–b)], the diagrams indicate that achieving either a sandwich state [Fig. 7(a)] or a full reversal [Fig.7(b)] of the initial liquid phase requires higher diffusion as the droplet radius decreases. At this value of $\epsilon$, hopping in non-preferred directions remains appreciable, meaning that only larger droplets can persist when $D$ is small; smaller droplets dissipate over time. As a result, for $D\leq 0.5$, no droplets survive, regardless of their size (up to the maximum tested $r_{d}=10$). Notably, in Fig.7(b), the full reversal state observed at $D\geq 0.7$ and larger $r_{d}$ $(r_{d}\gtrsim 7)$ is marked by orthogonal motion of clusters associated with the droplet and the initial liquid, as detailed in AppendixE.
For larger $\epsilon$ [Fig. 7(c–d)], the suppressed hopping in non-preferred directions allows droplets with smaller radii to persist, making both the sandwich state and complete reversal achievable for higher diffusion values $(D\geq 0.6)$. However, at lower diffusion levels $(0.3\leq D\leq 0.5)$—indicated by blue diamonds—a droplet (in state $\sigma_{d}=1$ or $\sigma_{d}=2$) fully disrupts the original liquid phase (state $\sigma=3$), giving rise to a dense lane consisting of particles in its own spin state. Over time, however, this lane becomes unstable due to the spontaneous emergence of another transversely moving spin state, $\sigma_{d}\pm 1[\mathrm{mod}4]$, within it. This newly nucleated spin state then displaces the existing lane and forms a new one. The process iterates, with each newly established lane being replaced by another, driven by spontaneous nucleation of transversely polarized droplets. This dynamic cycling of lane creation and annihilation, governed by competing transverse spin states, has also been observed in Refs. jd2024MIP; TSAIM. An illustration of this behavior for $D=0.3$ is provided in movie06 at Ref. zenodo.